# Moduli spaces of spacefilling branes in symplectic 4-manifolds

### v1

@@ 1 Introduction
This paper is concerned with a specific geometric structure defined on a symplectic manifold $(M,\omega)$, namely spacefilling branes structures. They are provided by closed $2$-forms $F$ such that the endomorphism
@@
squares to $-Id_{TM}$. In this case, $I$ is an integrable complex structure, and $F+i\omega$ is a holomorphic symplectic form (in particular, $\omega$ is of type [^1] This is in contrast to Kähler forms, which are real of type $(1,1)$. $(2,0)+(0,2)$ w.r.t. $I$, and $\dim_{\mathbb{R}}(M)$ is a multiple of $4$). Conversely, on the symplectic manifold $(M,\omega)$, a complex structure $I$ can be used to construct a holomorphic symplectic form – and thus a spacefilling brane – if and only if $F(\cdot,\cdot):=\omega(I\cdot,\cdot)$ is a 2-form, i.e. antisymmetric [^2] This is in contrast with Kähler geometry, where $\omega(I\cdot,\cdot)$ is symmetric and positive definite.. The space of spacefilling branes for a fixed symplectic form $\omega$ can thus be viewed as the space of precisely those holomorphic symplectic forms with fixed imaginary part $\omega$.
@@
We will restrict ourselves to the symplectic setting, and no knowledge of generalized complex geometry is assumed. Higher-dimensional coisotropic branes in symplectic manifolds are much less studied than the lowest-dimensional branes, Lagrangian submanifolds. We are interested in the deformation theory of coisotropic branes: one reason is that it provides new examples of branes, another is that it closely related to the deformation theory of coisotropic submanifolds, which is rich and highly non-linear [16].
@@
Let $M$ be either $T^{4}$ or the K3 manifold, and fix a symplectic form $\omega$. Our main result is a statement about the moduli space $\mathcal{M}_{\omega}$, the quotient of the space of spacefilling branes on $(M,\omega)$ by the group $\operatorname{Symp}_{*}(M,\omega)$ of symplectomorphisms inducing the identity in cohomology. To state the result, consider the map
@@
The moduli space of spacefilling brane structures $\mathcal{M}_{\omega}$ is a non-Hausdorff smooth manifold, and non-compact. For $M$ a K3 surface (respectively the 4-torus), endowed with any symplectic form $\omega$, the moduli space $\mathcal{M}_{\omega}$ is of dimension $20$ (respectively $4$).
@@
The above result relies in an essential way on a version of the Local Torelli Theorem. Indeed, we encode spacefilling branes as complex structures $I$ such that $\omega(I\cdot,\cdot)$ is skew-symmetric. On a compact $4$-manifold admitting a complex structure which has trivial canonical bundle, there is a well-known way to encode any complex structure by a line of closed complex 2-forms. The Local Torelli Theorem as stated in [6], which holds under the Kähler condition and we recall in Theorem 3.7, is a statement about the induced map from the moduli space of complex structures to the projectivization of $H^{2}(M,\mathbb{C})$. A suitable restriction of this map yields the map $\Phi$ in (1).
@@
Acknowledgements: We thank Marco Gualtieri for crucial input about the local Torelli theorem he gave us at the beginning of this project, and Luka Zwaan for pointing us to the existence of smooth universal deformations for the proof of Lemma 4.9.
@@ 2 Spacefilling branes in symplectic manifolds
In this section we review spacefilling branes in symplectic manifolds, establishing several equivalent characterizations. The general notion of brane arises naturally in the context of both string theory and generalized complex geometry [10, §3]. Specialized to the symplectic context, it reads as follows:
@@
([4], Example 7.8) A coisotropic A-brane in the symplectic manifold $(M,\omega)$ is a pair $(Y,F)$ of a submanifold $Y\subset M$ and a closed 2-form $F\in\Omega^{2}(Y,\mathbb{R})$ s.t.
@@
It follows that $I$ is a transverse complex structure, i.e. on any submanifold transverse to $T^{\omega}_{Y}$ it is integrable. The fact that $I$ is a transverse complex structure is shown in [11, §4], and we will provide a direct computational argument below.
@@
The systematic study of coisotropic A-branes in symplectic manifolds is originally due to Kapustin-Orlov in [11], who establish that there are higher-dimensional branes supported on some coisotropic submanifolds. The description as in Definition 2.1 can be found in [11, §4].
@@
In this case, $\omega^{-1}\circ F$ defines a complex structure on $M$, with respect to which $F+i\omega$ is holomorphic symplectic. We will show this in Proposition 2.6; to this aim, we state and prove the following known result (see [11, §4] for one direction of this equivalence):
@@
Then $I$ is a complex structure, i.e. integrable, if and only if $\operatorname{d\!}F=0$.
@@
Thus we have the following characterization [^3] A similar characterization, in which $\mathcal{L}$ is replaced by the Levi Civita connection of a Riemannian metric preserved by $I$, can be found e.g. in [15, Lemma 5.4]. of the vanishing of $N_{I}$:
@@ 3. HS(3)
We recall a well-known correspondence, which we will use in §3.2. In order to be self-contained, we provide a proof. As mentioned in [6, Remark 0.1], for K3 surfaces this description appears to be due to Andreotti.
@@ 3.2 Kähler 4-manifolds and the local Torelli theorem
Consider first an oriented compact 4-manifold $M$ which admits a complex structure and a holomorphic symplectic form (equivalently, by Remark 3.3, so that the canonical bundle $K$ is trivial). Recall that Proposition 3.4 provides a correspondence between complex structures $I^{\prime}$ defining the given orientation and complex lines $\mathbb{C}\Omega^{\prime}$ in $\Omega^{2}(M,\mathbb{C})$ satisfying certain conditions. This correspondence induces a well-defined map
@@
called the period map, which to the class of a complex structure $I^{\prime}$ associates the line $\mathbb{C}[\Omega^{\prime}]$, where $\Omega^{\prime}$ is a holomorphic symplectic form w.r.t. $I^{\prime}$.
Here
@@
The local Torelli problem – the question of when a period map from equivalence classes of complex structures to cohomology is a local isomorphism – is a classical problem that has been studied for different classes of manifolds. The provenance of the result for K3 surfaces specifically appears to be somewhat unclear: In [6], Huybrechts explains that this result was mentioned by André Weil in [20] without attribution, but may originally be due to Andreotti. More recently, the problem has been explored in greater generality: For example, [13, Theorem 1’] establishes cohomological criteria for the local Torelli theorem to be satisfied and [8, Proposition 3.14] proves it for generalized Calabi-Yau manifolds. The results in both texts certainly apply to complex tori.
By Theorem 3.7, locally near $I$, the moduli space of complex structures in these cases is a complex manifold contained in the projectivization of the 2nd cohomology. In particular, changing the holomorphic symplectic form within its cohomology class results in an isomorphic complex structure. It is well-know that the domain of the period map $\mathcal{P}$ is non-Hausdorff.
For both 4-tori and K3 surfaces – i.e. the manifolds considered in Theorem 3.7 – there are also known results about the global properties of the period map. For the purpose of this text, we point out the following (for the K3 surface see [6, Thm. 0.4][3, §2.1], for the 4-torus we refer to the beginning of [3, §2.3]):
@@
4 Deformations of spacefilling branes on symplectic $4$-manifolds
This section contains our main results. Consider a compact symplectic 4-manifold admitting a spacefilling brane such that the complex structure is Kähler. This restricts the underlying smooth manifold to the 4-torus and the K3 manifold, as we have established in Remark 3.6. Our results are geometric statements about the moduli space of spacefilling brane structures: the moduli space is a non-Hausdorff smooth manifold, it is non-compact, of dimension $4$ for the 4-torus and $20$ for the K3 manifold, and it carries a natural Lorentzian metric. These results are presented in §4.3 and §4.4. The previous subsections §4.1 and §4.2 are devoted to a characterization of spacefilling branes on general symplectic 4-manifolds, and to the definition of moduli space of spacefilling branes.
@@ 4.1 Spacefilling branes on symplectic 4-manifolds
In this subsection we restrict the correspondence recalled in §3.1 by fixing the imaginary part of the complex 2-form. In the next proposition we recover the characterization of spacefilling branes in symplectic 4-manifolds first observed by Kapustin-Orlov [11, §3].
@@ 4.2 The moduli space of spacefilling branes
where $\operatorname{Diff}_{*}(M)$ was defined in (6).
@@
Our reason for the above choice of equivalence is that by applying the Local Torelli Theorem 3.7 in the next subsections we will obtain results about the moduli space $\mathcal{M}_{\omega}$, rather than the about the moduli space arising from $\operatorname{Symp}_{0}(M,\omega)$.
@@ 4.3.1 Smoothness of the restricted period map
To show the claimed existence of the representative of $A\in H^{2}(M,\mathbb{C})$, we argue as follows: There exists a smooth universal family of deformations $X\to S$ of $(M,I)$, parametrized by an open neighborhood $S$ of $\mathbb{C}[\Omega]$ in the period domain $\mathcal{Q}$. For K3 surfaces this follows from [9, §6.2.3, Corollary 2.7 and Proposition 2.8], and for 4-tori from [14, Thm. VII.1, Thm. I. 52, Ex. I.54]. Denote by $I^{\prime}$ the complex structure on the fiber over $\mathbb{C}[\Omega^{\prime}]$; then $I^{\prime}$ lies nearby $I$, and $I^{\prime}$ is a representative of the class of complex structures that corresponds to $\mathbb{C}[\Omega^{\prime}]$ under Theorem 3.7. Let $\Omega^{\prime}$ be the unique 2-form lying in the complex line of forms associated to $I^{\prime}$ (as in Proposition 3.4) and whose imaginary part in cohomology is $[\omega]$ (the uniqueness is ensured by Lemma 4.8). Then $\Omega^{\prime}$ is a representative of $A$ lying nearby $\Omega=F+i\omega$ and satisfying HS(1)-HS(3). In particular, writing $\Omega^{\prime}=F^{\prime}+i\omega^{\prime}$, we have that $\omega^{\prime}$ lies nearby $\omega$ and lies in the same cohomology class, so by Moser’s theorem there is $\varphi\in\operatorname{Diff}_{*}(M)$ such that $\varphi^{*}\omega^{\prime}=\omega$. The desired representative of $A$ is $\varphi^{*}(\Omega^{\prime})$. ∎
@@ 4.4 Global properties of the moduli space of branes on Kähler $4$-manifolds
For $M$ a K3 manifold (respectively the 4-torus), endowed with any symplectic form $\omega$, the moduli space of brane structures $\mathcal{M}_{\omega}$ is not compact, and of real dimension $20$ (respectively $4$).

### v2

@@ 1 Introduction
This paper is concerned with a specific geometric structure defined on a symplectic manifold $(M,\omega)$, namely spacefilling brane structures. They are provided by closed $2$-forms $F$ such that the endomorphism
@@
squares to $-Id$. In this case, $I$ is an integrable complex structure, and $F+i\omega$ is a holomorphic symplectic form (in particular, $\omega$ is of type [^1] This is in contrast to Kähler forms, which are real of type $(1,1)$. $(2,0)+(0,2)$ w.r.t. $I$, and $\dim_{\mathbb{R}}(M)$ is a multiple of $4$). Conversely, on the symplectic manifold $(M,\omega)$, a complex structure $I$ can be used to construct a holomorphic symplectic form – and thus a spacefilling brane – if and only if $F(\cdot,\cdot):=\omega(I\cdot,\cdot)$ is a 2-form, i.e. antisymmetric [^2] This is in contrast with Kähler geometry, where $\omega(I\cdot,\cdot)$ is symmetric and positive definite.. The space of spacefilling branes for a fixed symplectic form $\omega$ can thus be viewed as the space of precisely those holomorphic symplectic forms with fixed imaginary part $\omega$.
@@
We will restrict ourselves to the symplectic setting, and no knowledge of generalized complex geometry is assumed. Higher-dimensional coisotropic branes in symplectic manifolds are much less studied than the lowest-dimensional branes, Lagrangian submanifolds. We are interested in the deformation theory of coisotropic branes: one reason is that it provides new examples of branes, another is that it is closely related to the deformation theory of coisotropic submanifolds, which is rich and highly non-linear [18].
@@
Let $M$ be either $T^{4}$ or the K3 manifold, and fix a symplectic form $\omega$ admitting a spacefilling brane. Our main result is a statement about the moduli space $\mathcal{M}_{\omega}$, the quotient of the space of spacefilling branes on $(M,\omega)$ by the group $\operatorname{Symp}_{*}(M,\omega)$ of symplectomorphisms inducing the identity in cohomology. To state the result, consider the map
@@
Let $M$ be the K3 manifold or the 4-torus, endowed with any symplectic form $\omega$ admitting a spacefilling brane structure.
The moduli space of spacefilling brane structures $\mathcal{M}_{\omega}$ is a non-Hausdorff smooth manifold, and non-compact. If $M$ is the K3 manifold, the moduli space $\mathcal{M}_{\omega}$ is of dimension $20$; if $M$ is the 4-torus, the moduli space is of dimension $4$.
@@
The above result relies in an essential way on a version of the Local Torelli Theorem, as we now outline.
First we encode spacefilling branes as those complex structures $I$ such that $\omega(I\cdot,\cdot)$ is skew-symmetric, via the assignment $F\mapsto I:=\omega^{-1}\circ F$, see Proposition 2.6 ii).
On a compact $4$-manifold admitting a complex structure which has trivial canonical bundle, there is a well-known way to encode any complex structure by a line of closed complex 2-forms, which we recall in Proposition 3.4. Complex structures satisying the above skew-symmetry property correspond to complex 2-forms with imaginary part $\omega$.
The Local Torelli Theorem as stated in [7], which holds under the Kähler condition and we recall in Theorem 3.7, is a statement about the induced map from the moduli space of complex structures to the projectivization of $H^{2}(M,\mathbb{C})$. A suitable restriction of this map yields the map $\Phi$ in (1).
@@
Acknowledgements: We thank Marco Gualtieri for crucial input about the local Torelli theorem he gave us at the beginning of this project, and Luka Zwaan for pointing us to the existence of smooth universal deformations for the proof of Lemma 4.9. We also thank Ignacio Barros for interesting discussions, Jim Stasheff for feedback, and the referee for valuable suggestions that improved the manuscript.
@@ 2 Spacefilling branes in symplectic manifolds
In this section we review spacefilling branes in symplectic manifolds, establishing several equivalent characterizations. The general notion of brane arises naturally in the context of both string theory and generalized complex geometry [12, §3]. Specialized to the symplectic context, it reads as follows:
@@
([5], Example 7.8) A coisotropic A-brane in the symplectic manifold $(M,\omega)$ is a pair $(Y,F)$ of a submanifold $i_{Y}\colon Y\hookrightarrow M$ and a closed 2-form $F\in\Omega^{2}(Y,\mathbb{R})$ s.t.
@@
It follows that $I$ is a transverse complex structure, i.e. on any submanifold transverse to $T^{\omega}_{Y}$ it is integrable. The fact that $I$ is a transverse complex structure is shown in [13, §4], and we will provide a direct computational argument below.
@@
The systematic study of coisotropic A-branes in symplectic manifolds is originally due to Kapustin-Orlov in [13], who establish that there are higher-dimensional branes supported on some coisotropic submanifolds. The description as in Definition 2.1 can be found in [13, §4].
@@
In this case, $\omega^{-1}\circ F$ defines a complex structure on $M$, with respect to which $F+i\omega$ is holomorphic symplectic. We will show this in Proposition 2.6; to this aim, we state and prove the following known result (see [13, §4] for one direction of this equivalence):
@@
Then the following holds: $I$ is a complex structure, i.e. integrable, if and only if $\operatorname{d\!}F=0$.
@@
Thus we have the following characterization [^3] A similar characterization, in which $\mathcal{L}$ is replaced by the Levi Civita connection of a Riemannian metric preserved by $I$, can be found e.g. in [17, Lemma 5.4]. of the vanishing of $N_{I}$:
@@ 3. HS(3)
We recall a well-known correspondence, which we will use in §3.2. In order to be self-contained, we provide a proof. As mentioned in [7, Remark 0.1], for K3 surfaces this description appears to be due to Andreotti.
@@ 3.2 Kähler 4-manifolds and the local Torelli theorem
Consider first an oriented compact 4-manifold $M$ which admits a complex structure and a holomorphic symplectic form (equivalently, by Remark 3.3, so that the canonical bundle $K$ is trivial). We denote by $\mathbf{Q}$ the subset of $H^{2}(M,\mathbb{C})$, cut out by the conditions
Recall that Proposition 3.4 provides a correspondence between complex structures $I^{\prime}$ defining the given orientation and complex lines $\mathbb{C}\Omega^{\prime}$ in $\Omega^{2}(M,\mathbb{C})$ satisfying certain conditions. This correspondence induces a well-defined map
@@
called the period map, which to the class of a complex structure $I^{\prime}$ associates the line $\mathbb{C}[\Omega^{\prime}]$, where $\Omega^{\prime}$ is a holomorphic symplectic form w.r.t. $I^{\prime}$. Here
@@
The local Torelli problem – the question of when a period map from equivalence classes of complex structures to cohomology is a local isomorphism – is a classical problem that has been studied for different classes of manifolds. The provenance of the result for K3 surfaces specifically ([7, page 4][11, Proposition 2.8 in Ch. 6]) appears to be somewhat unclear: In [7], Huybrechts explains that this result was mentioned by André Weil in [22] without attribution, but may originally be due to Andreotti. More recently, the problem has been explored in greater generality: For example, [15, Theorem 1’] establishes cohomological criteria for the local Torelli theorem to be satisfied and [9, Proposition 3.14] proves it for generalized Calabi-Yau manifolds. The results in both texts certainly apply to complex tori.
By Theorem 3.7, locally near $I$, the moduli space of complex structures in these cases is a complex manifold contained in the projectivization of the 2nd cohomology. In particular, changing the holomorphic symplectic form within its cohomology class results in an isomorphic complex structure. It is well-known that the domain of the period map $\mathcal{P}$ is non-Hausdorff.
For both 4-tori and K3 surfaces – i.e. the manifolds considered in Theorem 3.7 – there are also known results about the global properties of the period map. For the purpose of this text, we point out the following (for the K3 surface see [7, Thm. 0.4][4, §2.1], for the 4-torus we refer to the beginning of [4, §2.3]):
@@
Remark 3.10.
A version of the Local Torelli theorem holds for compact hyperkähler manifolds in general. ([1], see also [10, Theorem 4.1 and §3.1]). In this paper we will not address the question of whether our main result - Theorem 4.15, which relies on the Local Torelli Theorem - can be extended to compact hyperkähler manifolds of higher dimensions.
4 Moduli spaces of spacefilling branes on symplectic $4$-manifolds
This section contains our main results. Consider a compact symplectic 4-manifold admitting a spacefilling brane such that the complex structure is Kähler. Then this complex structure admits a holomorphic symplectic form, by Proposition 2.6. This restricts the underlying smooth manifold to the 4-torus and the K3 manifold, as we have established in Remark 3.6. Our results are geometric statements about the moduli space of spacefilling brane structures: the moduli space is a non-Hausdorff smooth manifold, it is non-compact, of dimension $4$ for the 4-torus and $20$ for the K3 manifold, and it carries a natural Lorentzian metric. These results are presented in §4.3 and §4.4. The previous subsections §4.1 and §4.2 are devoted to a characterization of spacefilling branes on general symplectic 4-manifolds, and to the definition of moduli space of spacefilling branes.
@@ 4.1 Spacefilling branes on symplectic 4-manifolds
In this subsection we restrict the correspondence recalled in §3.1 by fixing the imaginary part of the complex 2-form. In the next proposition we recover the characterization of spacefilling branes in symplectic 4-manifolds first observed by Kapustin-Orlov [13, §3].
@@ 4.2 The moduli space of spacefilling branes
where $\operatorname{Diff}_{*}(M)$ was defined in (7). Our reason for this choice of equivalence relation is based on the Local Torelli Theorem: by applying the Local Torelli Theorem 3.7 in the next subsections we will obtain results about the moduli space $\mathcal{M}_{\omega}$, rather than about the moduli space arising from $\operatorname{Symp}_{0}(M,\omega)$.
@@ 4.3.1 Smoothness of the restricted period map
To show the claimed existence of the representative of $A\in H^{2}(M,\mathbb{C})$, we argue as follows: There exists a smooth universal family of deformations $X\to S$ of $(M,I)$, parametrized by an open neighborhood $S$ of $\mathbb{C}[\Omega]$ in the period domain $\mathcal{Q}$. For K3 surfaces this follows from [11, §6.2.3, Corollary 2.7 and Proposition 2.8], and for 4-tori from [16, Thm. VII.1, Thm. I. 52, Ex. I.54]. Denote by $I^{\prime}$ the complex structure on the fiber over $\mathbb{C}[\Omega^{\prime}]$; then $I^{\prime}$ lies nearby $I$, and $I^{\prime}$ is a representative of the class of complex structures that corresponds to $\mathbb{C}[\Omega^{\prime}]$ under Theorem 3.7. Let $\Omega^{\prime}$ be the unique 2-form lying in the complex line of forms associated to $I^{\prime}$ (as in Proposition 3.4) and whose imaginary part in cohomology is $[\omega]$ (the uniqueness is ensured by Lemma 4.8). Then $\Omega^{\prime}$ is a representative of $A$ lying nearby $\Omega=F+i\omega$ and satisfying HS(1)-HS(3). In particular, writing $\Omega^{\prime}=F^{\prime}+i\omega^{\prime}$, we have that $\omega^{\prime}$ lies nearby $\omega$ and lies in the same cohomology class, so by Moser’s theorem there is $\varphi\in\operatorname{Diff}_{*}(M)$ such that $\varphi^{*}\omega^{\prime}=\omega$. The desired representative of $A$ is $\varphi^{*}(\Omega^{\prime})$. ∎
@@ 4.3.2 Smoothness of the moduli space of branes
Recall that the smooth manifold $M$ in this theorem must be either the K3 manifold or the 4-torus, by Proposition 2.6 and Remark 3.6.
@@ 4.4 Global properties of the moduli space of branes on Kähler $4$-manifolds
For $M$ a K3 manifold (respectively the 4-torus), endowed with any symplectic form $\omega$ admitting a spacefilling brane, the moduli space of brane structures $\mathcal{M}_{\omega}$ is not compact, and of real dimension $20$ (respectively $4$).